What is the standard deviation of the returns on a stock given the following information?

State of

Economy

Probability of

State of Economy

Rate of Return

if State Occurs

Boom

.28

.175

Normal

.67

.128

Recession

.05

.026

Group of answer choices

3.42 percent

4.01 percent

3.89 percent

3.28 percent

3.57 percent
Answer #1
State of Economy   Probability of state of Economy  

Rate of return if state occurs  

Boom 0.28 0.175
Normal 0.67 0.128
Recession 0.05 0.026

Expected rate of return when the probabilities (pi) of state of Economies are given is calculated using the formula:

Expected Return = E[R] = p1*R1 + p2*R2 + p3*R3 = 0.28*0.175 + 0.67*0.128 + 0.05*0.026 = 0.049 + 0.08576 + 0.0013 = 0.13606

Expected Return = E[R] = 0.13606

Now, Varicance is calculated using the formula:

Variance = σ2 = p1*(R1-E[R])2 + p2*(R2-E[R])2 + p3*(R3-E[R])2

Varince = σ2 = 0.28*(0.175-0.13606)2 + 0.67*(0.128-0.13606)2 + 0.05*(0.026-0.13606)2 = 0.000424570608 + 0.0000435256119999998 + 0.00060566018 = 0.0010737564

Standard Deviation is square-root of variance

Standard Deviation = σ = (0.0010737564)1/2 = 0.0327682224113546 = 3.27682224113546% ~ 3.28% (Rounded to two-decimals)

Answer -> 3.28 percent